OpenStudy (anonymous):

How many different ways can the first 10 letters of the alphabet be arranged? 4,628,800 3,628,800 2,628,800 1,628,800

OpenStudy (anonymous):

10!=3628800 but u better learn the basics first

OpenStudy (anonymous):

A pizza place offers 4 different cheeses and 10 different toppings. In how many ways can a pizza be made with 1 cheese and 4 toppings? 850 840 830 820

OpenStudy (anonymous):

its 830?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

820.

OpenStudy (anonymous):

4C1*10C4 = 4*210 = 820 ways

OpenStudy (anonymous):

no.. u can choose the cheese in 4C1=4 ways, topping in 10C4=210 ways so, total 210*4=840 ways

OpenStudy (anonymous):

btw, post separate questions in the question slot